# Estimating the link budget of satellite-based Quantum Key Distribution (QKD) for uplink transmission through the atmosphere

### v1

@@ 4 Quantum signal and final key rate
4.1 QKD utilizing a weak coherent pulse source
In this case, a weak coherent state serves as the source employed in the QKD experiment. The photon count of each pulse follows a Poisson distribution, where Alice’s expectation for the photon count is determined by the parameter $\mu$. Due to the statistical nature of photon numbers, some laser pulses may contain more than one photon, making them susceptible to an attack known as photon number splitting by an adversary. Once Bob discloses the basis of his measurement, Eve isolates a single photon from the pulse and preserves it in a quantum memory for subsequent measurement. Subsequently, Eve employs the same methodology to measure the stored photon, obtaining complete knowledge about multiphoton pulses without detection  [45].
The decoy pulse technique was devised as a countermeasure to thwart attacks. In this approach, Alice adjusts the average photon number $\mu$ of randomly spaced pulses known as decoys, which are not used in generating the secure key. By employing decoy pulses, stricter limitations are imposed on the amount of information Eve can gather from multiphoton transmissions. This is because Eve cannot determine in advance whether a given pulse is a signal or a decoy, making it more challenging for her to extract useful data.
@@ 4.1 QKD utilizing a weak coherent pulse source
The asymptotic key rate per laser pulse has a lower constraint using the decoy pulse approach which is
@@
In this context, the parameters are defined as follows: $q=1/2$ represents the basis reconciliation factor, $Q_{\mu}$ denotes the signal gain, which is the average photon number ratio of Bob’s detections to Alice’s pulses $\mu$. The quantum bit error rate for signal pulses is represented as $E_{\mu}$.
Furthermore, $f(E_{\mu})=1.22$ stands for the error correction efficiency for practical error correction codes. The binary entropy function is denoted as $H_{2}(x)$. Finally, $Q_{1}$ and $E_{1}$ correspond to the estimated gain and error rate for single-photon pulses. where $N_{\mu}$ $(N_{\nu})$ is the number of Bob’s received signal (decoy) counts. The security parameter $\Delta$ is given by
@@
$N$ is the length of the raw key, where $\epsilon$ is the total allowable probability that the final key is insecure, choose to be $\epsilon=10^{-9}$,$\epsilon_{EC}=10^{-10}$,
4.2 QKD using an entangled photon source
A model of a spontaneous parametric down-conversion (SPDC) source is employed to generate correlated photons. These entangled photons are then shared between Alice and Bob to be used for the QKD experiment[46]. When utilizing a perfect single photon source as the input, the key rate can be expressed as follows:
$K=\frac{1}{2}\Big{[}1-H_{2}(E+\xi)-f(E)H_{2}(E)-\Delta/N\Big{]}\,,$ (24)
Here $E+\xi$ is the modified QBER of the signal source. The quantity $E$ is the actual QBER in the asymptotic limit which can be computed directly by utilizing experimental results. The quantity $\xi$ is the statistical deviation of the asymptotic QBER due to finite measurement outcomes $N$. The quantity $\xi$ is a function of two parameters, the total number of signals sent by Alice $N$ and the Hilbert space dimension of the state sent by Alice. The term $f(E)H_{2}(E)$ is essentially equal to the measure of the key bits that are transmitted over the public communication channel for error correction purposes. The quantity $\Delta$ is the number of key bits required for error correction and privacy amplification. $H(x)$ is the binary entropy function.
Fig. 3 displays the final key rate for both Weak Coherent Pulse (WCP) and entangled sources. The X-axis represents the source rate of the transmitter, while the Y-axis represents the final key rate. The graph includes four detectors, indicating the ratio of the detection window to the repetition period (1.0 ns to 10 ns), and Alice’s detection efficiency of 0.25.
Figure 3: Final key rate for WCP and entangled source
@@ 4.2 QKD using an entangled photon source
Observatories
Loss (dB)
WCP (bits/sec)
Entangled (bits/sec)
IAO Hanle
44.28
2079.55
1052.11
ARIES Nainital
48.81
857.25
99.23
Mount-Abu
47.50
1009.16
107.56
Table 5: Link budget analysis for 3 different locations in India
From Table 5, it can be concluded that WCP sources have higher key rates compared to entangled sources.
@@ 5 Conclusion
We have presented simulation results analyzing the link performance at three different locations in India. The source being at the ground station, atmospheric effects become more prominent. Consequently, we compared atmospheric absorption using MODTRAN and calculated turbulence-related losses for all observatories. Our findings reveal that the Indian Astronomical Observatory (IAO) in Hanle exhibits the best performance, with total losses of approximately 44 dB. It’s important to note that due to turbulence, the beam also experiences increased divergence. Therefore, we have opted for an optimal aperture size of 15 cm, as going smaller results in higher losses, while going larger leads to a more turbulent beam, which in turn reduces the Strehl ratio at the detector. To further enhance performance, additional technologies like Adaptive Optics will be employed.
Declarations
The authors declare no competing interests. SRB performed the analysis under US’s supervision. Both authors contributed to the writing of the manuscript. We acknowledge the Indian Space Research Organisation for their support towards funding this research through the QuEST (Quantum Experiments using Satellite Technology) research grant. We acknowledge Animesh Sinha Roy for his technical assistance with the quantum signal and final key rate analysis. SRB thanks Saumya Ranjan Behera, Melvee George, and Mehak Layal for proofreading the manuscript. Lastly, we thank Kallon Sen for his technical assistance.
@@ Appendix A Quantum communication in Canary Island:
From [14], Table A1 gives the total specifications in detail for the two islands’ communication.
@@
Fried parameter $r_{0}$ 5cm (500 nm)
Table 6: Parameter used to calculate loss between two islands
@@
The following Table A2 presents the measured parameters during a specific nighttime observation. The absorption- and scattering-induced losses, denoted as $L_{A}$, for light at wavelengths 532 nm and 850 nm appear to range between 6 dB and 18 dB and 10 dB and 19 dB, respectively.
@@
Table 7: Different types of losses between two different wavelengths of two islands
@@ A.1 Validation of losses observed in Canary Island:
The MODTRAN parameter is given in Table A3 for the plot of atmospheric transmittance with the wavelength for the Canary Island  [14] is given in Fig. A1. From this plot, it can be concluded that the transmittance at 850 nm matches the loss given in Table A2.
@@
Table 8: Parameter used for MODTRAN plot for Canary islands
Figure 4: Atmospheric transmittance of Canary island
@@ A.2 Verification of Canadian data through simulation:
The Canadian Quantum Encryption and Science Satellite mission at the University of Waterloo proposed the uplink as well as downlink QKD experiment for the Ottawa location in Canada for an LEO orbit satellite. This location is situated in the sub-Arctic region with an altitude of 70m, and a sea-level rural atmosphere with a visibility of 5 km  [13]. The plot is given in Fig. A2.
Figure 5: Atmospheric transmittance of Canadian location
@@ A.3 Modtran parameters for three Indian observatories:
Table A4 gives the Modtran parameter used for three Indian observatories. Fig.A3 gives a comparison atmospheric transmittance plot for these three observatories.
Atmospheric parameter IAO Aries Mountabu
$H_{2}O$ (g/cm3) 0.0865 0.2489 0.1732
@@
Table 9: Atmospheric parameter consideration MODTRAN plot for IAO Hanle, Aries Nainital and Mount-Abu
Figure 6: Atmospheric transmittance comparison of three observatories
@@ A.4 Telescope size optimization:
The atmosphere plays a major role when a beam in the optical domain propagates through it. Turbulence degrades its quality. The turbulence strength can be measured with Fried parameter $r_{0}$. The main objective is to have a bigger value of the Fried parameter. If D= diameter of the beam from the telescope then a quantity $\frac{D}{r_{0}}$ is a ratio of the diameter of the beam from the telescope to the diameter of the turbulence strength i.e. Fried parameter. The lesser the value is better. Ideally, the value is $\leq 1$. For $(\frac{D}{r_{0}})>1$ signifies turbulated beam at the focus point which degrades the quality of the beam as shown in Fig 7. However, after $(\frac{D}{r_{0}})>2$ the degradation of beam starts rapidly  [26]. It can be concluded that $(\frac{D}{r_{0}})$ in between 1 to 2 is safe to operate.
@@
Which implies that $S$ decreases with increasing $(\frac{D}{r_{0}})$ causing difficult to detect at the detector  [48].
@@
$\langle r_{c}^{2}\rangle=0.73Z\sec\zeta\left(\frac{\lambda}{2W_{0}}\right)\left(\frac{2W_{0}}{r_{0}}\right)^{5/6}$ (27)
Where $\langle r_{c}^{2}\rangle$ is variance of beam shifting and $sec(\zeta)$ is the elevation angle of the beam from the ground station.
@@ A.4.1 Beam divergence under turbulence:
$\langle I(x,y,t)\rangle=\frac{P_{t}G_{a}}{4\pi Z^{2}}\eta_{t}\eta_{r}\eta_{atm}I_{u}\exp\left(-\frac{4(x^{2}+y^{2})}{\Phi_{B}^{2}Z^{2}}\right)$ (28)
  • $\eta_{t,r,atm}$ represents the transmittance of the transmitter, receiver telescopes, and atmosphere respectively.
@@
$\sigma_{jit}^{2}$ 1-axis 1- jitter variance which can be represented by
@@
Figure 7: Beam wander and Strehl ratio losses at different beam size
From Fig. A4. it can be concluded that the beam size 14-15 cm has the minimum losses. Table A4 gives the final loss calculated for IAO Hanle for different beam sizes.
@@
Loss in dB 47.47 46.57 46.33 46.22 46.12 46.04 45.7
Beam wander loss 3.53 1.27 0.27 0.12 0 0 0
Strehl ratio loss 0 1.17 1.89 2.20 2.56 3.79 4.86
Final loss in dB 51.0 49.0 48.49 48.52 48.68 49.83 50.56
Table 10: Final loss for different beam size

### v2

@@ 4.1 Effect of atmosphere on polarization degree of freedom
4.2 Simulated results from QKD utilizing a weak coherent pulse (WCP) source
QKD can be categorized into two primary types: Discrete Variable QKD (DV-QKD) and Continuous Variable QKD (CV-QKD). DV-QKD systems are further divided into prepare-and-measure protocols and entanglement-based protocols. In a typical prepare-and-measure protocol, Alice encodes each classical bit to a state of a quantum system and sends it to Bob, who performs a specific set of measurements on the incoming signals to retrieve the classical data encoded in their states. In the BB84 protocol, the most widely used prepare-and-measure QKD protocol, each classical bit is encoded into the polarization of a single photon. Alice randomly selects a basis between the horizontal/vertical (Z basis: H, V) and the +45°/-45° (X basis: A and D) and assigns bit values of 0 and 1 to these states. She then sends the chosen photon state to Bob. Bob randomly selects a basis to measure the incoming photon and records the result as a classical bit. After many transmissions, Alice and Bob publicly announce their chosen bases for each photon and discard cases where their bases do not match. They then estimate the error rate from a random subset of their data. If the error rate is low enough, below the security threshold, they apply error correction and privacy amplification to obtain a final shared secret key. A block diagram of BB84 using WCP is given in Fig. 3.
Figure 3: Block diagram of the BB84 QKD protocol using WCP source with source rate of 10 MHz.
[…]
In the practical implementation of weak coherent pulses in QKD, the quantum link becomes prone towards eavesdropping strategies like photon number splitting attacks. Hence, decoy states are introduced in many experimental demonstrations to prevent potential Eaves from gaining any useful information during the communication [49].
In this work, we have assumed such a decoy state-based system and estimated key rate and QBER considering quantum links from various ground stations in India with fixed source parameters. The quantum bit error rate for a decoy-state protocol is represented as $E_{\mu}$ and given as [50],
$E_{\mu}=\frac{\frac{Y_{0}}{2}+e_{detector}(1-e^{(-\eta\mu)})}{Q_{\mu}}$ (22)
where $Y_{0}$ is the dark count of the detector, $Q_{\mu}$ denotes the signal gain, $\eta$ is the detection efficiency and $e_{detector}$ is the probability of erroneous detection.
@@ 4.2 Simulated results from QKD utilizing a weak coherent pulse (WCP) source
Following the same reference, the asymptotic key rate per pulse was calculated, which has a lower constraint using the decoy state approach. The key rate including the finite key rate analysis for the decoy state-based BB84 QKD protocol is given by
@@
In this context, the parameters are defined as follows: $q=1/2$ represents the basis reconciliation factor, $f(E_{\mu})=1.22$ stands for the error correction efficiency for practical error correction codes. The binary entropy function is denoted as $H_{2}(x)$. Finally, $Q_{1}$ and $E_{1}$ correspond to the estimated gain and error rate for single-photon pulses. where $N_{\mu}$ $(N_{\nu})$ is the number of Bob’s received signal (decoy) counts. The security parameter $\Delta$ is given by
@@
where $N$ is the length of the raw key,
The QBER and key rate calculations for three different observatories are given in Table.5
Observatories Loss (dB) QBER for WCP in % Keyrates for WCP in bps
IAO Hanle 35.91 2.29 122.48
ARIES Nainital 37.78 2.96 73.55
Mount-Abu 37.19 2.72 86.63
Table 5: Estimated QBER and keyrate for 3 different locations in India, considering WCP source-based decoy state QKD protocol. The key rate calculated for the asymptotic limit as given in [20].
It can be concluded that all three observatories have QBER less than the information-theoretic security threshold of 11% meant for BB84 protocol.
4.3 Results from QKD simulations using an Entangled photon source
Entanglement-based protocols differ from prepare-and-measure systems by eliminating the need for active state encoding. In this work, we consider the BBM92 QKD protocol. Alice and Bob share a source of maximally entangled photon pairs. These photon pairs are entangled in polarization, with one photon belonging to Alice and the other to Bob. Both Alice and Bob independently choose to measure the photons on either a linear or diagonal polarization basis[51]. A block diagram of an Entangled source-based QKD protocol is given in Fig. 4.
The transmitter module includes an entangled photon source of 10 MHz pair generation rate. Both Alice and Bob, acting as receivers, have identical polarization-measuring optical setups, each followed by single-photon detectors. A block diagram of BBM92 using entangled photon source is given in Fig.4
Figure 4: Block diagram of the BBM92 QKD protocol using entangled photon source with 10MHz pair generation rate.
The asymptotic key rate per signal for the BBM92 QKD protocol is given by
$K=\frac{1}{2}Q_{\lambda}\Big{[}1-H_{2}(E)-f(E)H_{2}(E)\Big{]}\,,$ (25)
Here, $E$ is the QBER of the signal source. The total number of signals sent by Alice $N$. The term $f(E)H_{2}(E)$ is essentially equal to the measure of the key bits that are transmitted over the public communication channel for error correction purposes. $H(x)$ is the binary entropy function. All the parameters used in the equations have been taken from [52]. The QBER and key rates calculations for Entangled sources are given in Table.6 for the three observatories.
Observatories Loss (dB) QBER for Entangled source in % Keyrates for Entangled source in bps
IAO Hanle 35.91 5.62 14.08
ARIES Nainital 37.78 5.78 8.75
Mount-Abu 37.19 5.72 10.2
Table 6: QBER and key rate calculation for 3 different locations in India for Entangled sources.
It can be concluded that all three observatories have QBER less than the information-theoretic security threshold of 11% meant for the BBM92 protocol.
Fig. 5 and Fig. 6 display the final key rate for both Weak Coherent Pulse (WCP) and Entangled sources for all 3 observatories. The X-axis represents the source rate of the transmitter, while the Y-axis represents the final key rate. The graph includes four detectors, indicating the ratio of the detection window to the repetition period (1.0 ns to 10 ns), and Alice’s detection efficiency of 0.25.
Figure 5: Key rate estimation with BB84 protocol considering WCP source for three different observatories.
Figure 6: Key rate estimation considering an Entangled source-based QKD for the three potential observatory locations.
@@ 4.3 Results from QKD simulations using an Entangled photon source
Observatories Loss (dB) WCP (bits/sec) Entangled (bits/sec)
IAO Hanle 35.91 122.48 14.08
ARIES Nainital 37.78 73.55 8.75
Mount-Abu 37.19 86.63 10.2
Table 7: Link budget analysis for 3 different potential locations as ground stations in India.
From Table.7, it can be concluded that WCP sources have higher key rates compared to Entangled sources at all the three observatories. QBER for both types of sources for all three observatories is less than 11%. The schematic provided for the QKD experiments confirms these results.
@@ 5 Conclusion
We have presented simulation results analyzing the link performance at three different locations in India. The source being at the ground station, atmospheric effects become more prominent. Consequently, we compared atmospheric absorption using MODTRAN and calculated turbulence-related losses for all observatories. Our findings reveal that the Indian Astronomical Observatory (IAO) in Hanle exhibits the best performance, with total losses of approximately 36 dB. It’s important to note that due to turbulence, the beam also experiences increased divergence. Therefore, we have opted for an optimal aperture size of 15 cm, as going smaller results in higher losses, while going larger leads to a more turbulent beam, which in turn reduces the Strehl ratio at the detector. To further enhance performance, additional technologies like Adaptive Optics will be employed.
6 Declarations
6.1 Ethical Approval and Consent to participate
The authors US and SRB have followed the research ethics guidelines during the entire course of the work. The work has been completed with the full consent to participation from both the contributing authors.
6.2 Consent for publication
All the contributing authors (US and SRB) consent to the publication of this work in the journal. Both authors contributed to the writing of the manuscript. No other authors are involved.
6.3 Availability of supporting data
The data generated during the project has been included in the manuscript. The data pertaining to this work is not available online. Further data can be provided by the authors on request.
6.4 Competing interests/Authors’ contributions
The authors of this work are not aware of any competing interests or conflicts regarding this work. The authors thank Animesh Sinha Roy for his technical assistance with the quantum signal and final key rate analysis. SRB thanks Saumya Ranjan Behera, Melvee George, and Mehak Layal for proofreading the manuscript. The authors thank Kallol Sen for his technical assistance.
6.5 Funding
The work has been provided by the Indian Space Research Organization (ISRO) through the QUEST (Quantum Experiments using Satellite Technology) research grant.
@@ Appendix A Quantum communication in Canary Island:
From [14], Table 8 gives the total specifications in detail for the two islands’ communication.
@@
Fried parameter $r_{0}$ 5cm (at 500 nm wavelength)
Table 8: Parameters used to calculate loss between La Palma and Tenerife island.
@@
The following Table 9 presents the measured parameters during a specific nighttime observation. The absorption- and scattering-induced losses, denoted as $L_{A}$, for light at wavelengths 532 nm and 850 nm appear to range between 6 dB and 18 dB and 10 dB and 19 dB, respectively.
@@
Table 9: Different types of losses for different wavelengths between the two considered islands.
@@ A.1 Validation of losses observed in Canary Island:
The MODTRAN parameter is given in Table 10 for the plot of atmospheric transmittance with the wavelength for the Canary Island  [14] is given in Fig. 7. From this plot, it can be concluded that the transmittance at 850 nm matches the loss given in Table 9, since it is within the range given in the ref [14].
@@
Table 10: Parameters used in MODTRAN simulations for Canary Islands.
Figure 7: Atmospheric transmittance of Canary islands. X-axis represents wavelength in nm from 500 to 1500. Y-axis represents the transmittance value.
@@ A.2 Verification of Canadian data through simulation:
The Canadian Quantum Encryption and Science Satellite mission at the University of Waterloo proposed the uplink as well as downlink QKD experiment for the Ottawa location in Canada for an LEO orbit satellite. This location is situated in the sub-Arctic region with an altitude of 70m, and a sea-level rural atmosphere with a visibility of 5 km  [13]. The plot is given in Fig. 8.
Figure 8: Atmospheric transmittance from the considered Canadian location(Ottawa).
@@ A.3 Modtran parameters for three Indian observatories:
Table 11 gives the Modtran parameter used for three Indian observatories. Fig.9 gives a comparison atmospheric transmittance plot for these three observatories.
Atmospheric parameter IAO Aries Mount-Abu
$H_{2}O$ ($g/cm^{2}$) 0.0865 0.2489 0.1732
@@
Table 11: Atmospheric parameters considered in MODTRAN simulations for IAO Hanle, Aries Nainital and Mount-Abu.
Figure 9: Atmospheric transmittance comparison of three potential observatory locations.
@@ A.4 Telescope size optimization:
The atmosphere plays a major role when a beam in the optical domain propagates through it. Turbulence degrades its quality. The turbulence strength can be measured with Fried parameter $r_{0}$. The main objective is to have a bigger value of the Fried parameter. If D= diameter of the beam from the telescope then a quantity $\frac{D}{r_{0}}$ is a ratio of the diameter of the beam from the telescope to the diameter of the turbulence strength i.e. Fried parameter. The lesser the value is better. Ideally, the value is 1$\leq$ 1≤ 1. For $(\frac{D}{r_{0}})>1$ indicates that the beam is more turbulent at the focal point, leading to a degradation in beam quality. As this value increases, it results in greater losses at the detector, as illustrated in the Fig.10. However, after $(\frac{D}{r_{0}})>2$ the degradation of beam starts rapidly  [26]. It can be concluded that $(\frac{D}{r_{0}})$ in between 1 to 2 is safe to operate.
Figure 10: The plot of increasing Strehl ratio loss in dB vs $\frac{D}{r_{0}}$ value.
@@
Which implies that $S$ decreases with increasing $(\frac{D}{r_{0}})$ causing difficulty in detection  [54]. To calculate attenuation due to turbulence, we use an approximation of the above as given by [31].
@@
$\sqrt{\langle r_{c}^{2}\rangle}=0.73Z\sec\zeta\left(\frac{\lambda}{2W_{0}}\right)\left(\frac{2W_{0}}{r_{0}}\right)^{5/6}$ (28)
Where $\langle r_{c}^{2}\rangle$ is variance of beam shifting and $sec(\zeta)$ is the elevation angle of the beam from the ground station. Further, the angular beam wander can be defined as $\theta_{BW}=\frac{\sqrt{\langle r_{c}^{2}\rangle}}{L}$, using which we calculate the loss due to beam wander in our considered uplink scenario [56].
@@ A.4.1 Beam divergence under turbulence:
$\langle I(x,y,t)\rangle=\frac{P_{t}G_{a}}{4\pi Z^{2}}\eta_{t}\eta_{r}\eta_{atm}I_{u}\exp\left(-\frac{4(x^{2}+y^{2})}{\Theta_{B}^{2}Z^{2}}\right)$ (29)
  • $\eta_{t,r,atm}$ represents the transmittance of the transmitter, receiver telescopes, and atmosphere, respectively.
@@
$\sigma_{jit}^{2}$ 1-axis 1- jitter variance, which can be represented by
@@
Figure 11: Beam wander and Turbulence losses for different initial beam sizes.
From Fig. 11. it can be concluded that the beam size 14-15 cm has the minimum losses. Table A5 gives the final loss calculated for IAO Hanle for different beam sizes.
@@
Beam wander loss 2.38 0.87 0.52 0.40 0.30 0.08 0.02
Turbulence loss 0.24 0.52 0.69 0.78 0.89 1.32 1.78
Total loss dependent on D 2.62 1.39 1.22 1.18 1.19 1.4 1.80
Table 12: Total loss due to turbulence added for different initial beam sizes.